OpenStudy (anonymous):

How does 5 log 3 +log 4=log972??

OpenStudy (rulnick):

4 times 3^5 = 972 log (4(3^5)) = log 4 + log 3^5 = log 4 + 5 log 3

OpenStudy (anonymous):

5log 3 +log 4= \[\log 3^{5}+ \log 4\] \[\log( 4\times 3^{5})\] log 972